OpenStudy (anonymous):

how do i find the missing diameter of a circle when the radius is given (6th grade math)

OpenStudy (tkhunny):

"Missing"? Where could it have gone? \(2\pi Radius = \pi Diameter = Circumference\)

OpenStudy (anonymous):

what about for a circle 6th grade math

OpenStudy (tkhunny):

It makes no difference. It is the same in Antarctica and on the moon and in PhD school and in Kindergarten. You may be assigned a value for \(\pi\). Maybe 3.14?